Suppose that there are no variable costs. Think about what this means. Marginal cost is the additional cost of producing one more unit of a good.

1) If there are only fixed costs, then what is the additional cost of producing an additional unit?
2) If a monopolist is running a business with only fixed costs what is price elasticity of demand at the profit maximizing level of output?
3) show that all monopolists facing positive marginal cost produce where demand is elastic?
4) Suppose that many small identical firms decide to join and make one large firm. Assume that this joining does not affect their costs. Graphically show how level of output and price differ between the many small firms and the single large firm acting as a monopoly. Show how consumer and producer surplus in the competitive industry differ from consumer and producer surplus in the monopoly industry. Why do economists believe that a monopoly industry is inefficient?
5) Which concept related to the long run average cost curve gives a large monopoly an advantage over several smaller firms?
Answer #1

P 22 dc q Foc 13p = 0 - 959 ) o -) paq paa + 2 2 2 2 = 0 - P+q 1 » aq e = -1 =- ana ar q than zero ap 29 . +2 ap > -) > -) or